OpenStudy (anonymous):

Add simplify your answer. 3x+8 4 ________ + _________ x^2-16 x^2-16

OpenStudy (saifoo.khan):

\[\frac{3x+8 + 4}{x^2 - 16} \to \frac{3x+12}{x^2 - 16}\to \frac{3(x+4)}{(x-4)(x+4)} \to solve.\]

OpenStudy (anonymous):

I don't understand your answer saifoo.khan, sorry.

OpenStudy (saifoo.khan):

No worries! Where are you stuck?

OpenStudy (anonymous):

How did you combine the two?

OpenStudy (anonymous):

Did you just bring over the 4?

OpenStudy (saifoo.khan):

same as, 1/2 + 3/2 We look for common denominators. (1+3)/2 4/2

OpenStudy (saifoo.khan):

which 4?

OpenStudy (anonymous):

3x+8 4 <--- this 4 ________ + _________ x^2-16 x^2-16

OpenStudy (saifoo.khan):

that's in the question, isn't it?

OpenStudy (anonymous):

Yes, I think I understand this now. Let me re look it.

OpenStudy (saifoo.khan):

Sure. dont forget to hit good answer.

OpenStudy (anonymous):

Wait hold one, where did the 12 and 16 go?

OpenStudy (anonymous):

on* not one

OpenStudy (saifoo.khan):

3x+12 can be written as, 3(x+4) -taking 3 common and x^2-16 can be written as (x-4)(x+4)

OpenStudy (anonymous):

Oh ok. Thank you, you helped me a lot. :D

OpenStudy (saifoo.khan):

welcome. :D